9/5log(base2)x=10.

Answer by stanbon(75887)

(9/5)log(base2)x=10
------------------
log2(x) = 10(5/9)
log2(x) = 50/9
---
x = 2^(50/9)
x = 47.03315....
===============
Cheers,
Stan H.